[Hallc_running] [New Logentry] RC Update

smithg at jlab.org smithg at jlab.org
Sat Feb 22 14:00:01 EST 2020

Logentry Text:
No RC meeting today.

The target experts have been making small tweaks to optimize the target polarization and have been successful raising it from ~52% to 54.9% over the past 24 h or so, even while 30 uA have been on the target. That's great! Then today, Mingyu rotated to transverse and we are taking 5 h of data there before going to longitudinal again.

Last night around 9 PM we had a lot of vertical beam motion, which the crew chief picked up on since they had been asked to keep an eye on this. He noticed some correlations, and made some changes (see his HClogs at 10 PM (https://logbooks.jlab.org/entry/3789212) & 10:20 PM (https://logbooks.jlab.org/entry/3789222)). Since then our BPMs have been looking pretty stable. See https://logbooks.jlab.org/entry/3789419 and https://logbooks.jlab.org/entry/3789604. Still a few small excursions (0.3-0.4 mm) but no big ones. It's too soon to say if the larger excursions we've seen are finally behind us, but so far, so good!

The PD reported this morning that the 4h access for injector PSS work Tuesday morning is firming up, which will likely be our (only) opportunity for an access next week to pump on the xfer line vacuum spaces in the spectrometers, as well as the Moller magnet. Let the RC know of any other access requests. 

Looking good so far this weekend!


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3789632
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200222/e9b3fca8/attachment.html>

More information about the Hallc_running mailing list